OpenStudy (anonymous):

when im finding the area for anything and i get a number like 12.4789 do i write all the decimal numbers?

OpenStudy (phi):

It depends. Sometimes you want the answer to the nearest (for example) hundred th, in which case you round If you are asked the area of a circle, often people do not multiply out pi in pi r^2 for example if r=2 the area is 4\(\pi\), and rather than giving a bunch of decimals they leave the answer 4\(\pi\)

OpenStudy (anonymous):

ok thanks